# EvoSeed: Unveiling the Threat on Deep Neural Networks with Real-World Illusions

### v1

@@ Abstract.
Deep neural networks are exploited using natural adversarial samples, which have no impact on human perception but are misclassified. Current approaches often rely on the white-box nature of deep neural networks to generate these adversarial samples or alter the distribution of adversarial samples compared to training distribution. To alleviate the limitations of current approaches, we propose EvoSeed, a novel evolutionary strategy-based search algorithmic framework to generate natural adversarial samples. Our EvoSeed framework uses auxiliary Diffusion and Classifier models to operate in a model-agnostic black-box setting. We employ CMA-ES to optimize the search for an adversarial seed vector, which, when processed by the Conditional Diffusion Model, results in an unrestricted natural adversarial sample misclassified by the Classifier Model. Experiments show that generated adversarial images are of high image quality and are transferable to different classifiers. Our approach demonstrates promise in enhancing the quality of adversarial samples using evolutionary algorithms. We hope our research opens new avenues to enhance the robustness of deep neural networks in real-world scenarios.
Deep Learning, Computer Vision, CMA-ES, Diffusion Models, Natural Adversarial Samples
@@
Project Website can be accessed at https://shashankkotyan.github.io/EvoSeed
1. Introduction
Figure 1. Illustration of EvoSeed’s performance in finding adversarial samples under various conditions compared to Random Search (RandSeed). Our proposed EvoSeed algorithm finds high-quality Natural Adversarial Samples with a high attack success rate.
Figure 2. Illustration of the EvoSeed framework to optimize initial seed vector to generate Unconstrained adversarial samples. In this approach to generating adversarial examples, a specific class from the Cifar10 dataset is chosen as the target for the attack, denoted by condition $c$. The process begins with a random initial seed vector $z$, optimized within an $L_{\infty}$ constraint using the Covariance Matrix Adaptation Evolution Strategy (CMA-ES). A conditional diffusion model $G$ utilizes the provided condition $c$ to iteratively reduce noise from the perturbed seed vector $z^{\prime}$, resulting in image distributions $x$. The CMA-ES’s state is updated based on the fitness of the perturbed seed vector $z^{\prime}$, determined by a classifier model $F$ and the condition $c$. This iterative process continues until an adversarial seed vector $z^{\prime}$ is found, which can be used to generate an adversarial sample.
@@ 1. Introduction
Deep Neural Networks have succeeded unprecedentedly in various visual recognition tasks. However, their performance decreases when the testing distribution differs from the training distribution, as shown by Hendrycks et al. (Hendrycks et al., 2021) and Ilyas et al. (Ilyas et al., 2019). This poses a significant challenge in developing robust deep neural networks capable of handling such shifts in distribution. Adversarial samples and adversarial attacks exploit this vulnerability by manipulating images to alter distribution compared to the original distribution.
@@
Research by Dalvi et al. (Dalvi et al., 2004) underscores that adversarial manipulations of input data often lead to incorrect predictions from classifiers, raising serious concerns about the security and integrity of classical machine learning algorithms. This concern remains relevant, especially considering that state-of-the-art deep neural networks are highly vulnerable to adversarial attacks involving deliberately crafted perturbations to the input (Madry et al., 2018; Kotyan and Vargas, 2022).
[…]
@@
While adversarial samples (Madry et al., 2018; Su et al., 2019; Kotyan and Vargas, 2022) expose vulnerabilities in deep neural networks, their artificial nature and reliance on constrained input data limit their real-world applicability. In contrast, the challenges become more pronounced in practical situations, where it becomes infeasible to include all potential threats within the training dataset comprehensively. This heightened complexity underscores the increased susceptibility of deep neural networks to Natural Adversarial Examples proposed by Hendrycks et al. (Hendrycks et al., 2021) and Unrestricted Adversarial Examples proposed by Song et al. (Song et al., 2018). These types of adversarial samples have gained prominence in recent years as a significant avenue in adversarial attack research, as they can make substantial alterations to images without significantly impacting human perception of their meanings and faithfulness.
@@
In this context, we present EvoSeed, the first Evolution Strategy-based algorithmic framework designed to genearate Unrestrcited Natural Adversarial Samples as shown in Figure 2. Our algorithm requires a Conditional Diffusion Model $G$ and a Classifier Model $F$ to generate adversarial samples $x$. Specifically, it leverages the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) at its core to enhance the search for adversarial seed vectors $z^{\prime}$ that can generate adversarial samples $x$. The CMA-ES fine-tunes the generation of adversarial samples through an iterative optimization process based on the Classification model outputs $F(x)$, utilizing them as fitness criteria for subsequent iterations. Ultimately, our objective is to search for an adversarial seed vector $z^{\prime}$ that, when used, causes our Conditional Diffusion Model $G$ to generate an adversarial sample $x$ misclassified by the Classifier Model $F$ and also is close to the training distribution of Conditional Diffusion Model $G$ and Classifier Model $F$, as shown in Figure 1.
@@
Framework to Generate Natural Adversarial Samples: We propose a model-agnostic black-box algorithmic framework based on an Evolutionary Strategy titled EvoSeed to generate unrestricted natural adversarial samples. Our framework can generate natural adversarial examples using any auxiliary conditional diffusion and classifier models, as shown in Figure 2.
@@
High-Quality Natural Adversarial Samples: Our results show that adversarial samples created using EvoSeed can also misclassify a variety of robust and non-robust classifiers alike, while being similar to the training distribution.
2. Related Work
Over the past few years, generative models such as Generative Adversarial Networks (Goodfellow et al., 2020) and Diffusion Models (Sohl-Dickstein et al., 2015) have emerged as leading tools for content creation and the precise generation of high-quality synthetic data. Several studies have employed creativity to generate Adversarial Samples; some propose the utilization of surrogate models such as (Xiao et al., 2018a; Chen et al., 2023b, a; Zhao et al., 2018; Lin et al., 2023; Jandial et al., 2019; Song et al., 2018), while other advocates the perturbation of latent representations as a mechanism for generating adversarial samples (Song et al., 2018; Zhao et al., 2018).
In the initial phases of devising natural adversarial samples, Xiao et al. (Xiao et al., 2018b) employs spatial warping transformations for their generation. Concurrently, Shamsabadi et al. (Shamsabadi et al., 2020) transforms the image into the LAB color space, producing adversarial samples imbued with natural coloration.
Song et al. (Song et al., 2018) proposes first to train an Auxiliary Classifier Generative Adversarial Network (AC-GAN) and then apply the gradient-based search to find adversarial samples under its model space. Another research proposes Adversarial GAN (AdvGan) (Xiao et al., 2018a), which removes the searching process and proposes a simple feed-forward network to generate adversarial perturbations and is further improved by Jandial et al. (Jandial et al., 2019). Similarly, Chen et al. (Chen et al., 2023b) proposes the AdvDiffuser model to add adversarial perturbation to generated images to create better adversarial samples with improved FID scores.
Yet, these approaches often have one or more limitations,
a) Some approaches rely on changing the distribution of generated images compared to the training distribution of the classifier, such as (Xiao et al., 2018b; Shamsabadi et al., 2020), or
b) Some approaches rely on the white-box nature of the classifier model to generate adversarial samples such as (Song et al., 2018; Chen et al., 2023b), or
c) Some approaches rely heavily on training models to create adversarial samples such as (Xiao et al., 2018a; Song et al., 2018; Jandial et al., 2019), or
d) Some approaches rely on generating adversarial samples for specific classifiers, such as (Xiao et al., 2018a; Jandial et al., 2019).
Thus, in contrast, we propose the EvoSeed algorithmic framework, which does not suffer from the limitations mentioned above in generating adversarial samples.
3. Background and Problem Formulation
The Diffusion Model is first proposed by Sohl-Dickstein et al. (Sohl-Dickstein et al., 2015) that can be described as a Markov chain with learned Gaussian transitions. It comprises of two primary elements: a) The forward diffusion process, and b) The reverse sampling process. The diffusion process transforms an actual distribution into a familiar straightforward random-normal distribution by incrementally introducing noise. Conversely, in the reverse sampling process, a trainable model is designed to diminish the Gaussian noise introduced by the diffusion process systematically.
Let us consider a true distribution represented as $x\in\mathbb{R}$, where $x$ can be any kind of distribution such as images (Ho et al., 2020; Dhariwal and Nichol, 2021; Ho et al., 2022; Ho and Salimans, 2022), audio (Kong et al., 2021; Huang et al., 2022a, b; Kim et al., 2022), or text (Li et al., 2022). The diffusion process is then defined as a fixed Markov chain where the approximate posterior $q$ introduces Gaussian noise to the data following a predefined schedule of variances, denoted as $\beta_{1},\beta_{2}\dots\beta_{T}$:
(1) $q(x_{1:T}|x_{0}):=\prod^{T}_{t=1}~{}q(x_{t}|x_{t-1})$
where $q(x_{t}|x_{t-1})$ is defined as,
(2) $q(x_{t}|x_{t-1}):=\mathcal{N}(x_{t};~{}\sqrt{1-\beta_{t}}\cdot x_{t-1},~{}\beta_{t}I).$
(3) $p_{\theta}(x_{0:t}):=p(x_{T})\cdot\prod^{T}_{t=1}~{}p_{\theta}(x_{t-1}|x),$
where $p_{\theta}(x_{t-1}|x)$ is defined as,
(4) $p_{\theta}(x_{t-1}|x_{t}):=\mathcal{N}\left(x_{t-1};~{}\mu_{\theta}(x_{t},t),~{}\Sigma_{\theta}(x_{t},t)\right).$
Furthermore, the generation process can be conditioned akin to various categories of generative models (Mirza and Osindero, 2014; Sohn et al., 2015). For instance, by integrating with text embedding models as an extra condition $c$, the conditional-based diffusion model $G_{\theta}(x_{t},c)$ creates content along the description (Ramesh et al., 2022; Saharia et al., 2022; Rombach et al., 2022; Nichol et al., 2022). This work mainly uses a conditional diffusion model to construct adversarial samples.
Unrestricted Adversarial Samples: We follow the definition from Song et al. (Song et al., 2018). Given that $\mathcal{I}$ represents a collection of images under consideration that can be categorized using one of the $K$ predefined labels. Let’s consider a testing classifier $f:\mathcal{I}\rightarrow\{1,2\dots K\}$ that can give a prediction for any image in $\mathcal{I}$. Similarly, we can consider an oracle classifier $o:O\subseteq\mathcal{I}\rightarrow\{1,2\dots K\}$ different from the testing classifier, where $O$ represents the distribution of images understood by the oracle classifier. An unrestricted adversarial sample can defined as any image inside the oracle’s domain $O$ but with a different output from the oracle classifier $o$ and testing classifier $f$. Formally defined as $x\in O$ such that $o(x)\neq f(x)$. The oracle $o$ is implicitly defined as a black box that gives ground-truth predictions. As it aligns with human assessment, the set $O$ should encompass all images perceived as realistic by humans.
@@ 3. Background and Problem Formulation
Optimization on Initial Seed Vector to Generate Adversarial Samples: Let’s define a conditional diffusion model $G$ that takes an initial seed vector $z$ and a condition $c$ to generate an image $x$. Based on this, we can define the image generated by the conditional diffusion model as,
(5) $x=G(z,c)\quad\text{where}\quad z\in\mathcal{N}(\mu,\alpha^{2})$
Here, $\mu$ and $\alpha$ depend on the chosen conditional model $G$.
@@
From the definition of the image classification problem, we can define a classifier $F$ such that $F(x)\in\mathbb{R}^{K}$ is the probabilities (confidence) for all the available $K$ labels for the image $x$. We can also define the soft label of a condition or confidence of the condition $c\in\{1,2\dots K\}$ as $F()_{c}$.
@@
(6) $z^{\prime}=z+\eta\quad\text{such that}\quad\text{argmax}~{}[F(~{}G(z+\eta,~{}c)~{})]\neq c$
@@
However, research by Poyuan et al. (Poyuan et al., 2023) reveals that the failure points are distributed everywhere inside the space, mostly generating images that cannot be associated with the condition $c$. To navigate these failure cases, we make the problem non-trivial by searching around the space of a well-defined initial random vector $z$. We do this by imposing an $L_{\infty}$ constraint on perturbation to initial seed vector $(\eta)$, so the modified problem becomes,

### v2

# Breaking Free: How to Hack Safety Guardrails in Black-Box Diffusion Models!

@@ Abstract
Deep neural networks can be exploited using natural adversarial samples, which do not impact human perception. Current approaches often rely on deep neural networks’ white-box nature to generate these adversarial samples or synthetically alter the distribution of adversarial samples compared to the training distribution. In contrast, we propose EvoSeed, a novel evolutionary strategy-based algorithmic framework for generating photo-realistic natural adversarial samples. Our EvoSeed framework uses auxiliary Conditional Diffusion and Classifier models to operate in a black-box setting. We employ CMA-ES to optimize the search for an initial seed vector, which, when processed by the Conditional Diffusion Model, results in the natural adversarial sample misclassified by the Classifier Model. Experiments show that generated adversarial images are of high image quality, raising concerns about generating harmful content bypassing safety classifiers. Our research opens new avenues to understanding the limitations of current safety mechanisms and the risk of plausible attacks against classifier systems using image generation.
CAUTION: This article includes model-generated content that may contain offensive or distressing material that is blurred and/or censored for publication.
@@
footnotetext: Project Website can be accessed at: https://shashankkotyan.github.io/EvoSeed
1 Introduction
Figure 1: Adversarial images created with EvoSeed are prime examples of how to deceive a range of classifiers tailored for various tasks. Note that, the generated natural adversarial images differ from non-adversarial ones, suggesting the adversarial images’ unrestricted nature.
@@ 1 Introduction
Deep Neural Networks have succeeded unprecedentedly in various visual recognition tasks. However, their performance decreases when the testing distribution differs from the training distribution, as shown by Hendrycks et al. [1] and Ilyas et al. [2]. This poses a significant challenge in developing robust deep neural networks capable of handling such shifts in distribution. Adversarial samples and adversarial attacks exploit this vulnerability by manipulating images to alter distribution compared to the original distribution.
@@
Research by Dalvi et al. [3] underscores that adversarial manipulations of input data often lead to incorrect predictions from classifiers, raising serious concerns about the security and integrity of classical machine learning algorithms. This concern remains relevant, especially considering that state-of-the-art deep neural networks are highly vulnerable to adversarial attacks involving deliberately crafted perturbations to the input [4, 5].
Various constraints are imposed on these perturbations, making these perturbations subtle and challenging to detect. For example, $L_{0}$ adversarial attack such as One-Pixel Attack [5, 6] limit the number of perturbed pixels, $L_{1}$ adversarial attack such as EAD [7] restrict the Manhattan distance from the original image, $L_{2}$ adversarial attack such as PGD-L2 [4] restrict the Euclidean distance from the original image, and $L_{\infty}$ adversarial attack such as PGD-L [4] restricts the amount of change in all pixels. Some of these attacks are of White-Box nature such as [4, 7], while others are of Black-Box nature such as [5, 6, 8]
@@
While adversarial samples [4, 5, 6] expose vulnerabilities in deep neural networks; their artificial nature and reliance on constrained input data limit their real-world applicability. In contrast, the challenges become more pronounced in practical situations, where it becomes infeasible to include all potential threats comprehensively within the training dataset. This heightened complexity underscores the increased susceptibility of deep neural networks to Natural Adversarial Examples proposed by Hendrycks et al. [1] and Unrestricted Adversarial Examples proposed by Song et al. [9]. These types of adversarial samples have gained prominence in recent years as a significant avenue in adversarial attack research, as they can make substantial alterations to images without significantly impacting human perception of their meanings and faithfulness.
@@
In this context, we present EvoSeed, the first Evolution Strategy-based algorithmic framework designed to generate Natural Adversarial Samples in an unrestricted setting as shown in Figure 2. Our algorithm requires a Conditional Diffusion Model $G$ and a Classifier Model $F$ to generate adversarial samples $x$ for a given classification task. Specifically, it leverages the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) at its core to enhance the search for adversarial initial seed vectors $z^{\prime}$ that can generate adversarial samples $x$. The CMA-ES fine-tunes the generation of adversarial samples through an iterative optimization process based on the Classification model outputs $F(x)$, utilizing them as fitness criteria for subsequent iterations. Ultimately, our objective is to search for an adversarial initial seed vector $z^{\prime}$ that, when used, causes our Conditional Diffusion Model $G$ to generate an adversarial sample $x$ misclassified by the Classifier Model $F$ and is also close to the human perception, as shown in Figure 1.
@@
Framework to Generate Natural Adversarial Samples: We propose a black-box algorithmic framework based on an Evolutionary Strategy titled EvoSeed to generate natural adversarial samples in an unrestricted setting. Our framework can generate adversarial examples for various tasks using any auxiliary conditional diffusion and classifier models, as shown in Figure 2.
@@
High-Quality Photo-Realistic Natural Adversarial Samples: Our results show that adversarial samples created using EvoSeed are photo-realistic and do not change the human perception of the generated image however can be misclassified by various robust and non-robust classifiers.
2 Optimization on Initial Seed Vector to Generate Adversarial Samples
Figure 2: Illustration of the EvoSeed framework to optimize initial seed vector $z$ to generate a natural adversarial sample. The Covariance Matrix Adaptation Evolution Strategy (CMA-ES) iteratively refines the initial seed vector $z$ and finds an adversarial initial seed vector $z^{\prime}$. This adversarial seed vector $z^{\prime}$ can then be utilized by the Conditional Diffusion Model $G$ to generate a natural adversarial sample $x$ capable of deceiving the Classifier Model $F$.
@@ 2 Optimization on Initial Seed Vector to Generate Adversarial Samples
Let’s define a Conditional Diffusion Model $G$ that takes an initial seed vector $z$ and a condition $c$ to generate an image $x$. Based on this, we can define the image generated by the conditional diffusion model $G$ as,
$x=G(z,c)\quad\text{where}\quad z\sim\mathcal{N}(\mu,\alpha^{2})$ (1)
here $\mu$ and $\alpha$ depend on the chosen Conditional Diffusion Model $G$.
@@
From the definition of the image classification task, we can define a classifier $F$ such that $F(x)\in\mathbb{R}^{K}$ is the probabilities (confidence) for all the available $K$ labels for the image $x$. We can also define the soft label or confidence of the condition $c\in\{1,2\dots K\}$ as $F(\cdot)_{c}$, where $\sum_{i=1}^{K}F(x)_{i}=1$.
@@
$z^{\prime}=z+\eta\quad\text{such that}\quad\operatorname*{arg\,max}~{}[F(~{}G(z+\eta,~{}c)~{})]\neq c$ (2)
@@
However, research by Poyuan et al. [10] reveals that the failure points are distributed everywhere inside the space, mostly generating images that cannot be associated with the condition $c$. To navigate these failure cases, we make the problem non-trivial by searching around the space of a well-defined initial random vector $z$. We do this by imposing an $L_{\infty}$ constraint on perturbation to initial seed vector $\eta$, so the modified problem becomes,